The warranty on a machine specifies that it will be replaced at failure or age 5 which ever occurs first. The machine's age at failure, X, has density function

f(x) = 1/7, 0 < x < 7

f(x) = 0 otherwise. Let Y be the age of the machine at the time of replacement. Calculate the expected value of Y.
Answer #1

E(Y) =E(min(x,5)) = min(x,5)f(x) dx = (x/7) dx + (5/7) dx

=(x2/14) |50 +(5x/7)|75

=(25/14)+(10/7)

=45/14

=3.21 years